High concentrations of ammonia (NH3), nitrite ion and nitrate ion in water can kill fish. Lethal concentration of these species for rainbow trout are 1.1 mg/L, 0.40 mg/L, and 1361 mg/L, respectively. Express these concentrations in molality units, assuming a solution density of 1.00 g/mL.
1
Expert's answer
2020-03-09T09:41:01-0400

Molecular weight of ammonia is 17.031 g/mol.

Molar concentration is defined by the following formula:

M = n / V

where M - molarity, n - number of moles, V - volume.

As n = m / Mr (m - mass, Mr - molecular weight, Mr(NH3) = 17.031 g/mol), the molar concentrations are:

M1 = 1.1 mg / 1 L = 1.1 g × 10-3 / (17.031 g/mol × 1 L) = 0.065 × 10-3 mol/L = 65 µmol/L = 65 µM

M2 = 0.40 mg / 1 L = 0.40 g × 10-3 / (17.031 g/mol × 1 L) = 0.024 × 10-3 mol/L = 24 µmol/L = 24 µM

M3 = 1361 mg / 1 L = 1361 g × 10-3 / (17.031 g/mol × 1 L) = 79.9 × 10-3 mol/L = 79.9 mmol/L = 79.9 mM


Answer: 65 µM, 24 µM, 79.9 mM.